OpenStudy (anonymous):

multiply 2 sqrt of x+1 and 2 sqrt of x-1

OpenStudy (lgbasallote):

\[2\sqrt{x+1} \times 2\sqrt{x-1} \implies (2\times 2) \times (\sqrt{x+1} \times \sqrt{x-1})\] does that help?

OpenStudy (anonymous):

am i not going to use the foil method?

OpenStudy (lgbasallote):

you're going to use the FOIL method too \[\implies 4 \times \sqrt{(x+1)(x-1)}\] you FOIL that (x+1)(x-1)

OpenStudy (anonymous):

but not in the first step of that ,thank you

OpenStudy (lgbasallote):

welcome

OpenStudy (lgbasallote):

and you use FOIL when you have something like \[(a+b) \times (c +d)\]this is more like \[ab \times cd\]

OpenStudy (anonymous):

but not the one like i gave

OpenStudy (lgbasallote):

lol nevermind my metaphor =))